# Two Shades of Quark Color: Parallel Canons across the Cold War Divide

### v1

@@ 4.3.1 Design choices that encode the Regge brief
Read through this lens, the E–36 chain displays the hallmarks of a theory-led, observables-first programme. The choice of an internal supersonic gas jet target technique—transferred from U–70 to the Main Ring—prioritized access to the strict forward limit for optical-theorem normalization and dispersive tests [66, 65, 64]. Correspondingly, the collaboration’s success metrics were amplitude-level quantities, $\sigma_{\mathrm{tot}}(s)$, $\rho(s)$, and the diffraction slope $B(s)$, analyzed in the grammar of diffractive models, analyticity, and Regge shrinkage rather than in a partonic language [62, 63, 14, 16, 15]. In this respect, Dubna’s intention to embed a senior theorist was not a matter of prestige but of methodological fit: the same S-matrix/Regge commitments that motivated the measurement programme also were expected to shape the interpretation.
That Fermilab ultimately hosted Matveev in the Theory Division rather than within the experimental roster did not alter the underlying dynamic: the experimental chain proceeded with theory-shaped choices of observables and tests characteristic of Dubna’s composite/analytic tradition, while remaining institutionally compatible with NAL’s clearer theorist–experimentalist boundary. The resulting publications on the forward slope and the $\rho$-parameter exemplify this division of labour: they test global amplitude constraints with high precision and without reliance on microscopic carrier degrees of freedom [62, 63, 16].
@@ 5 Discussion
The Dubna strand was embedded in a broadly holist S-matrix culture that prioritized constraints on observables—analyticity, unitarity, crossing, Regge behavior—over unobservable field variables. In this idiom, hadrons are organized by poles, residues, and trajectories in the complex-$J$ plane; explanation proceeds by global amplitude structure rather than by specifying fundamental carriers [15, 58, 57, 14, 16]. This observables-first stance resonates with the period’s composite models at Dubna and helps explain why “color” entered there as a statistical/structural attribute (a selection rule on residues and allowed couplings), not as a local gauge charge [5, 4, 11].
These commitments had concrete experimental consequences. In the joint JINR–Fermilab program, notably the E–36 internal-target chain, design choices favored direct access to forward elastic observables governed by S-matrix constraints: the use of a supersonic gas-jet target enabled optical-theorem normalization and precision tests of dispersion/Regge systematics [66, 65, 64]. The collaboration’s key publications reported the diffraction slope $B(s)$ and the ratio $\rho(s)$ at high energies—quantities that test analyticity, crossing, and Regge shrinkage rather than any particular partonic micro-mechanism [62, 63, 14, 16]. Institutionally, Dubna also expected senior theorists to lead experimental strategy and interpretation (an expectation visible around the E–36 era), whereas Fermilab maintained a sharper theorist/experimentalist division of labor [65]. In effect, the same S-matrix grammar that motivated the measurements supplied the success criteria against which E–36 was judged.
@@
Against this backdrop, it is unsurprising that B. V. Struminsky’s hidden quantum number and the Dubna composite framing, though relying on the same SU(3) triplicity that appears in Han–Nambu, did not couple to a gauge-field ontology—the cornerstone of QCD [5, 4, 3, 8]. The absence of Dubna’s 1965 move from many Western QCD narratives is therefore not best explained by politics alone; it also reflects a language gap between a structural, S-matrix program and a dynamical, gauge-theory program that, after 1974, became progressively dominant [16, 32]. On the account developed here, the Dubna color proposal belongs to a parallel, structuralist canon: successful and illuminating within its regime, but ultimately orthogonal to the reductionist path that furnished the decisive high-energy discriminators for color as a local gauge charge.
@@
Unlike reductionism, which grounds explanation in fundamental micro-entities, structural realism emphasizes inter-theoretic coherence and the stability of formal structure across scales, offering a framework where explanatory power resides in the architecture of relations, not in ontological primitives. Conversely, QCD embodies a refined or scale-sensitive reductionism, where micro-dynamical laws are seen as generative of macro-phenomena, but the actual practice and conceptual structure of the theory reveal certain holist and effective-theoretic elements. It supports reduction in principle, but in practice depends on autonomous models, structural constraints, and non-trivial emergence. Therefore, the dominance of QCD in the 1970s was less the result of a strategic convergence and more a product of the pragmatic alignment of theory and calculational tools—shaped by communal priorities and institutional dynamics. Subsequent developments suggest that a domain-sensitive pluralism, balancing structuralist and reductionist approaches, has been more productive for advancing physics. In line with epistemic pluralism and its practical implementation in physics, the post-positivist theoretical-operational model of complex experiments [73] offers an explicit operational description of arbitrary complex experiments in terms of both phenomenal and instrumental theories. It demonstrates their interdependence, since phenomenal theories inform the selection of preparation and measurement frameworks. Moreover, it shows that distinct theoretical groundings for the same phenomena entail different sequences of their operational deformations.
@@ Acknowledgments
The author is grateful to Valerie Higgins, Archivist and Historian at Fermilab, for her assistance in providing access to the E-36 chain files from the Fermilab Archives.
Endnote: Boris Vladimirovich Struminsky (1939–2003) was a Soviet, Russian, and Ukrainian theoretical physicist. From 1962 to 1965, he was a doctoral student at the Steklov Institute of Mathematics in Moscow, where he continued as a junior researcher in 1965–1966. From 1966 to 1971, Struminsky worked as a researcher at the Joint Institute for Nuclear Research (JINR) in Dubna. Beginning in 1971, he held research positions at the Institute for Theoretical Physics of the Academy of Sciences of the Ukrainian SSR—now the Bogolyubov Institute for Theoretical Physics of the National Academy of Sciences of Ukraine—in Kyiv. [74].

### v2

@@ 4.3.1 Design choices that encode the Regge brief
Read through this lens, the E–36 chain displays the hallmarks of a theory-led, observables-first programme. The choice of an internal supersonic gas jet target technique—transferred from U–70 to the Main Ring—prioritized access to the strict forward limit for optical-theorem normalization and dispersive tests [68, 67, 66]. Correspondingly, the collaboration’s success metrics were amplitude-level quantities, $\sigma_{\mathrm{tot}}(s)$, $\rho(s)$, and the diffraction slope $B(s)$, analyzed in the grammar of diffractive models, analyticity, and Regge shrinkage rather than in a partonic language [64, 65, 15, 17, 16]. In this respect, placing a senior theorist in the loop was less a prestige move than a methodological one—meant to align the measurement programme with solution-based (S-matrix/Regge) reasoning, while, given Matveev’s quark focus, also opening the door to quark-level modifications and interpretations.
That Fermilab ultimately hosted Matveev in the Theory Division rather than within the experimental roster did not alter the underlying dynamic: the experimental chain proceeded with theory-shaped choices of observables and tests characteristic of Dubna’s composite/analytic tradition, while remaining institutionally compatible with NAL’s clearer theorist–experimentalist boundary. The resulting publications on the forward slope and the $\rho$-parameter exemplify this division of labour: they test global amplitude requirements with high precision and without reliance on microscopic carrier degrees of freedom [64, 65, 17].
@@ 5 Discussion
The Dubna strand was embedded in a broadly holist S-matrix culture that prioritized requirements of observables—analyticity, unitarity, crossing, Regge behavior—over unobservable field variables. In this idiom, hadrons are organized by poles, residues, and trajectories in the complex-$J$ plane; explanation proceeds by global amplitude structure rather than by specifying fundamental carriers [16, 60, 59, 15, 17]. This observables-first stance resonates with the period’s composite models at Dubna and helps explain why “color” entered there as a statistical/structural attribute (a selection rule on residues and allowed couplings), not as a local gauge charge [5, 4, 12].
These commitments had concrete experimental consequences. In the joint JINR–Fermilab program, notably the E–36 internal-target chain, design choices favored direct access to forward elastic observables governed by S-matrix equations: the use of a supersonic gas-jet target enabled optical-theorem normalization and precision tests of dispersion/Regge systematics [68, 67, 66]. The collaboration’s key publications reported the diffraction slope $B(s)$ and the ratio $\rho(s)$ at high energies—quantities that test analyticity, crossing, and Regge shrinkage rather than any particular partonic micro-mechanism [64, 65, 15, 17]. Institutionally, Dubna also expected senior theorists to lead experimental strategy and interpretation (an expectation visible around the E–36 era), whereas Fermilab maintained a sharper theorist/experimentalist division of labor [67]. In effect, the same S-matrix grammar that motivated the measurements supplied the success criteria against which E–36 was judged.
@@
With this in mind, it is unsurprising that B. V. Struminsky’s hidden quantum number and the Dubna composite framing, though relying on the same SU(3) triplicity that appears in Han–Nambu, did not couple to a gauge-field ontology—the cornerstone of QCD [5, 4, 3, 9]. The absence of Dubna’s 1965 move from many Western QCD narratives is therefore not best explained by politics alone; it also reflects a language gap between a structural, S-matrix program and a dynamical, gauge-theory program that, after 1974, became progressively dominant [17, 34]. On the account developed here, the Dubna color proposal belongs to a parallel, structuralist canon: successful and illuminating within its regime, but ultimately orthogonal to the reductionist path that furnished the decisive high-energy discriminators for color as a local gauge charge.
@@
Unlike reductionism, which grounds explanation in fundamental micro-entities, structural realism emphasizes inter-theoretic coherence and the stability of formal structure across scales, offering a framework where explanatory power resides in the architecture of relations, not in ontological primitives. Conversely, QCD embodies a refined or scale-sensitive reductionism, where micro-dynamical laws are seen as generative of macro-phenomena, but the actual practice and conceptual structure of the theory reveal certain holist and effective-theoretic elements. It supports reduction in principle, but in practice depends on autonomous models, structural criteria, and non-trivial emergence. Therefore, the dominance of QCD in the 1970s was less the result of a strategic convergence and more a product of the pragmatic alignment of theory and calculational tools—shaped by communal priorities and institutional dynamics. Subsequent developments suggest that a domain-sensitive pluralism, balancing structuralist and reductionist approaches, has been more productive for advancing physics. In line with epistemic pluralism and its practical implementation in physics, the post-positivist theoretical-operational model of complex experiments [75] offers an explicit operational description of arbitrary complex experiments in terms of both phenomenal and instrumental theories. It demonstrates their interdependence, since phenomenal theories inform the selection of preparation and measurement frameworks. Moreover, it shows that distinct theoretical groundings for the same phenomena entail different sequences of their operational deformations.
@@ Acknowledgments
The author is grateful to Valerie Higgins, Archivist and Historian at Fermilab, for her assistance in providing access to the E-36 chain files from the Fermilab Archives. I thank Prof. V. A. Petrov (IHEP, Protvino) for stimulating discussions that sharpened several arguments presented here.
Endnotes